Grade 12Mechanics

The speed of a projectile at its greatest height, is 1.414/2.236 times its speed at half the maximum height. The angle of projection=?

Aoishi
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Dear Student,
As we know that,
The maximum heoght for the projectile is:
H = u sintheta/g
h/2 = 1.414/2.236 sinѲ/10
5h =1.414/2.236 sinѲ
5√5h/√2 = sinѲ
Ѳ = sin-1(5√5h/√2).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)